# Deep Bayes Factors

### v2

@@ 4 Generative AI for Bayesian Model Criticism
We conclude this section with further discussion of possible model criticism developments using generative Bayes in the context of existing literature. Gelfand et al. (1992) and Box (1980) argue that model criticism should be addressed using posterior predictive distributions which are consonant with the intended use of the model. In addition, predictive distributions are compatible across models while posteriors are not. The methodology of Gelfand et al. (1992) rests upon the ability to obtain desired predictive distributions and to calculate expectations under these distributions. These authors argue for a cross-validation viewpoint, comparing an observed value against conditional predictive distributions arising from single point deletions. In particular, one of their suggestions is to construct a discrepancy measure (checking function) comparing each observation $Y_{i}^{0}$ against a posterior expectation $E[Y_{i}^{0}\,|\>Y_{\backslash i}^{0}]$ where $Y_{\backslash i}^{0}$ are all but the $i^{th}$ observation inside $Y^{(n)}_{0}$. The conditional expectation after removing the observation $Y_{i}^{0}$ is not too easy to compute and they suggested importance sampling. We suggest one feasible computational strategy based on generative posterior samplers. Using the Bayesian GAN approach (Wang and Ročková, 2022), one can train a GAN mapping $g(Y_{n-1},X)$ to simulate from $\pi(\theta\,|\>Y_{n-1})$ by passing a random noise variable $X$ through this mapping for a fixed vector $Y_{n-1}$ of length $n-1$. This training is performed solely on fake data. One can then plug a vector $Y_{\backslash i}^{0}$ into $g(Y_{\backslash i}^{0},X)$ to perform iid sampling from the posterior distribution $\pi(\theta\,|\>Y_{\backslash i}^{0})$ for any $i=1,\dots,n$. This sampling is cost-free because it is merely based on passing noise through the trained mapping evaluated at $Y_{\backslash i}^{0}$. The ability to simulate from this posterior makes simulating from $\pi(Y_{i}\,|\>Y_{\backslash i}^{0})$ and, therefore, approximating $E[Y_{i}\,|\>Y_{\backslash i}^{0}]$ feasible. One can, for example, compute the discrepancy measure $Z(Y_{0})=\frac{1}{n}\sum_{i=1}^{n}(Y_{i}^{0}-E[Y_{i}\,|\>Y_{\backslash i}^{0}])^{2},$ which avoids the use of observation $Y_{i}^{0}$ for the predictive check on $Y_{i}^{0}$. A tail area evidence measure can be then obtained by assessing the location of $Z(Y_{0})$ relative to the a reference distribution of $Z(Y)$ where, for example, $Y$ has been simulated from the (posterior predictive under) model $M$.
@@ 5.1 Estimation Consistency
The strength of our generative approach to Bayes factor estimation is that we can evaluate the estimator at any data $Y$, not just the observed data $Y^{(n)}_{0}$. This allows us to recreate distributions of Bayes factor estimators when $Y$ is generated from either model $M_{1}$ and model $M_{2}$. Inspecting entire distributions of Bayes factors $BF_{1,2}(Y)$ provides a more comprehensive image of discriminability among the two models. The quality of our estimator can be gauged by the closeness of the distributions of true and estimated Bayes factors. Figure 5 portrays cartoons of plausible shapes of Bayes factor distributions (true and estimated) under the two models. The relative position of the true (or estimated) $BF_{1,2}(Y_{0})$ within the plot provides stronger information than merely deciding whether the Bayes factor is above or below $1$. In addition, we estimate the KL divergence between empirical distributions (kernel density estimates (KDE)). These distances are computed separately under each model and then subsequently weighted by the prior model probability. Similarly, we measure the Spearman’s rho correlation (rank-based) between the estimated and true BF’s for simulated data under each model and take their weighted (by the model prior probability) sum.
We also compute an unbiased estimator of the prior $\pi(M_{j})$ by averaging $\pi(M_{j}\,|\>Y)$ over the simulated data $(Y_{1},...,Y_{T_{0}})^{\prime}\sim P(Y)$, denoted as $\hat{\pi}(M_{j})=\frac{1}{T_{0}}\sum_{i=1}^{T_{0}}\pi(M_{j}\,|\>Y_{i}).$ If $\hat{\pi}(M_{j})$ deviates significantly from true $\pi(M_{j})$, the estimator may not be trusted Schad et al. (2022). In our numerical examples, we assume $\pi(M_{1})=\pi(M_{2})=0.5$.
@@ 5.2 Measures of Surprise
Our simulation-based framework lends itself naturally to inference with Bayes factors based on tail-area surprise measures. While the $p$-value is a tail area (as opposed to likelihood-based) evidence measure, it can be calibrated to obtain a lower bound to a Bayes factor (Bayarri and Berger, 1998). Before seeing the data, Bayes factors are random variables and we can think of them as a proxy for a test statistic. Evaluating the position of the Bayes factor $BF_{1,2}(Y_{0}^{(n)})$ within the distributions of $BF_{1,2}(Y)$ for $Y\sim\pi(\cdot\,|\>M_{1})$ and $Y\sim\pi(\cdot\,|\>M_{2})$ has a flavor of the prior predictive $p$-value popularized by Box (1980) . We consider surprise measures
@@
which are essentially prior predictive $p$-values using the Bayes factor as a departure statistic. We do not argue that the prior-predictive reference distribution is preferred over alternatives including the conditional predictive $p$-value (Bayarri and Berger, 1998) which is justifiable when there is suspicion that poor priors were used and which is provably superior to various Bayesian $p$-values in large samples (Robins et al., 2000). In simulation-based models considered here, however, it might be difficult to come up with appropriate statistics needed for the computation of the conditional predictive $p$-values. While our computational framework enables the use of posterior-predictive $p$-value (Guttman, 1967; Rubin, 1984) by simulating $Y$ from the posterior (not prior) predictive under each mode, we focus on the prior predictive $p$-values acknowledging its weaknesses in its prior sensitivity. Using surprise measures for Bayes factors as inferential tools is not a new concept. For example, Vlachos and Gelfand (2003) suggest a simulation-intensive approach to calculate various tail area measures of surprises for Bayes factors and García-Donato and Chen (2005) propose a Monte Carlo method to find a decision rule $c$ such that $P(B_{1,2}(Y)\leq c|M_{1})=P(B_{1,2}(Y)\geq c|M_{2})$. Computing such surprise values was also considered in Schönbrodt and Wagenmakers (2018) who proposes a Bayes factor design analysis aimed at determining an appropriate sample size $n$ that ensures that with a high probability the sample data from the models exhibit strong Bayes factors in favor of the true model. When closed-form evaluations of Bayes factors are unavailable, the computation of surprise measures can be involved. The ability to rapidly evaluate the Deep Bayes factor mapping for any $Y\in{\rm Supp}(P)$ facilitates the computation of surprise measures. For example, we approximate $p_{1}(\cdot)$ and $p_{2}(\cdot)$ in (13) by replacing $BF_{1,2}(\cdot)$ with an estimator $\widehat{BF}_{1,2}(\cdot)$ and by replacing tail probabilities with empirical proportions based on simulated data $Y$. We denote these approximations by $\hat{p}_{1}(\cdot)$ and $\hat{p}_{2}(\cdot)$. Since the inferential quality of a Bayes factor estimator can be assessed by how well it approximates the tail probabilities in (13), we measure the MSE between the actual surprise and the estimated surprise by $MSE_{surp}=\sum_{j=1}^{2}\pi(M_{j})~{}\hat{\mathbb{E}}_{Y\sim M_{j}}\left[p_{j}(Y)-\hat{p}_{j}(Y)\right]^{2}.$
Figure 5: Thick line: the true BF distributions on samples simulated from $M_{1}$ or $M_{2}$ (depending on the color); Colored area: the estimated BF distributions on samples simulated from $M_{1}$ or $M_{2}$ (depending on the color).
@@ 5.3 Model Inference
The actual precision in estimating Bayes factors may be less relevant for decision making when there is overwhelming evidence for one of the two models (and the estimation error is appropriately small). It is the quality of decisions with Bayes factor’s that ultimately matters, not the approximation quality of its estimator. Jeffreys’ Bayes factor scale (Jeffreys, 1998) is a suggested interpretation of the evidence strength based on the magnitude of the Bayes factor binned into 10 categories. The coarsest categorization is into two classes, based on the decision boundary at $1$ (values greater than one favor model $1$ and values smaller than one favor model $2$). To quantify the inference quality for various thresholds (not just $1$), we use the area under the ROC curve (AUC) computed as follows. Considering $1$ as a “positive” label and $2$ as “negative”, we can calculate the true positive rate (TPR) and the false positive rate (FPR). Given the decision threshold $c\in(0,\infty),$ we write TPR and FPR as $TPR_{c}(\widehat{BF}_{1,2})=\mathbb{P}(\widehat{BF}_{1,2}(Y)>c|Y\sim M_{1})$ and $FPR_{c}(\widehat{BF}_{1,2})=\mathbb{P}(\widehat{BF}_{1,2}(Y)>c|Y\sim M_{2})$, where $\mathbb{P}$ is the empirical distribution of a simulated sample from the model. As these values depend on the choice of $c$, we can draw the ROC curve representing the relationship between TPR and FPR when $c$ is sent from 0 to $\infty$. The AUC does not penalize biased or improperly scaled estimation.
@@ 6.1 Toy Examples
In this section, we focus on the Negative Binomial vs. Poisson distributions in Example 1. For the other two datasets, see Section S4. This dataset can represent various scenarios with different hyperparameters $(\alpha_{1},\beta_{1},\alpha_{2},\beta_{2})$, which determine the landscape of the decision boundary of the model selection problem. Here, we choose $(\alpha_{1},\beta_{1},\alpha_{2},\beta_{2})=(2,2,4,4)$, which we call Dataset 1 (i), and $(\alpha_{1},\beta_{1},\alpha_{2},\beta_{2})=(1,1,1,1)$ as in Example 1, which we call Dataset 1 (ii). The former is more separable near decision boundaries than the latter in the sense that the probability that $BF_{1,2}(Y)$ is near $1$ is smaller (see, Figure S21 in Section S5). The true Bayes factor is available in a closed form (Bernardo and Smith, 2009). Letting $Y_{\cdot}=\sum_{i=1}^{n}Y_{i}$,
@@
Because of the explicit likelihood, we can readily use the well-known Bridge sampling (Bennett, 1976; Meng and Wong, 1996) and Warp Bridge sampling (Meng and Schilling, 2002) as baselines. We set the number of MCMC iterations of Bridge and Warp sampling as 150,000 with the burn-in 10,000. We also evaluate the ABC-based Bayes factor estimator (Robert et al., 2011), which is commonly employed in applications without an explicit likelihood. To compute the ABC-based estimates for a large number of simulated samples within a reasonable timeframe, we adopt a rank-based rule for the ABC algorithm instead of the standard $\epsilon$-acceptance rule. This approach allows for the reuse of ABC samples across different simulated datasets simultaneously, significantly reducing computation time. The detailed algorithm is provided in Section S2, and hyperparameter choices are listed in Table S2 therein. Consequently, the top 120 ABC samples are used to compute the ABC-based estimator in (2) with a small adjustment to avoid numerical explosion. As a result, the range of the ABC-based estimators is approximately [1/120,120]. We vary the sample size from $n=2^{1}$ to $n=2^{7}$, specifically focusing on sample sizes under 1,000. Studying such small sample sizes remains relevant for practical purposes, e.g., in cognitive and psychological science, sample sizes often range from 30 to 60 (Schad et al., 2022) and a sample size of 180 is considered large (Schad et al., 2022). For each $Y_{0}^{(n)}$, we run our algorithm and the baseline methods 20 times to assess estimation stability. As discussed in Section S3, we use both FNNs and BNNs for our algorithm, which we will refer to as DeepBF-FNN and DeepBF-BNN, respectively, where BNNs stand for batch normalization neural networks.
@@
We begin by assessing estimation consistency, as illustrated in Figures 6 and S22 (Section S5) for a sample size $n=128$. The first row of these figures presents scatter plots that compare the estimated $\widehat{BF}_{1,2}(Y)$ with the true Bayes factors $BF_{1,2}(Y)$ from simulated samples $Y$. The scatter plot for DeepBF-BNN suggests unbiased and properly scaled estimates, while DeepBF-FNN exhibits more reserved scaling. In Figure S22, both Bridge and Warp methods yield well-scaled and unbiased estimates. In contrast, Figure 6 shows bias in both methods, although biased estimates may still hold utility within a rank-based decision rule, as reflected by the surprise measures in (13). The ABC method appears non-informative. Additional insights are offered in Figures S24 and S25 (Section S5), which present scatter plots for increasing values of $n$. Here, the ABC method provides more accurate estimations for smaller $n$ values, but these estimations become non-informative as $n$ increases. An increase in the number of ABC samples, currently at 1,200,000, could potentially improve these outcomes. On the contrary, Bridge and Warp show unstable estimates for $Y\sim M_{2}$ but improve as $n$ becomes larger. While DeepBF-FNN tends to outperform DeepBF-BNN, its efficacy wanes with larger $n$, possibly due to the network’s fixed intermediate width across all $n$ values to stabilize optimization (Section S3).
@@
In the second row of Figure 6 and S22 (Section S5), we have empirical distribution plots of the Bayes factors (true and estimated) under the two models. While the KDE shapes of these plots corroborate the scatter plot shapes, we can also see the separability of distributions under each model, and the overall distributional deviation of the estimates. We also visualize the estimation quality for an instance $Y_{0}^{(n)}$ drawn from $M_{2}$. In Figure 6 (f-j), each vertical red line represents the true $BF_{1,2}(Y_{0}^{(n)})$, and the green lines represent its estimates from 20 different trials. For DeepBF-BNN, the green lines cluster closely around the red line, whereas for all other methods, the green lines show more deviation. The blue line is a reference line at the decision threshold of 1. Our algorithm’s estimates tend to be more conservative, closer to the blue line, while the Warp and Bridge estimates are overly optimistic, diverging further from the blue line. The variance in the multiple estimates increases in the following order ABC $<$ Warp and Bridge $<$ DeepBF-BNN $<$ DeepBF-FNN. More comprehensively, the scatter plots among multiple DeepBF-BNN estimates are presented in Figure S26 (Section S5). The plots in Figure S22 (f-j) display similar results, but with Warp and Bridge estimates appearing unbiased and clustering tightly around the true red line.

### v3

@@ 4 Generative AI for Bayesian Model Criticism
We conclude this section with further discussion of possible model criticism developments using generative Bayes in the context of existing literature. [21] and [13] argue that model criticism should be addressed using posterior predictive distributions which are consonant with the intended use of the model. In addition, predictive distributions are compatible across models while posteriors are not. The methodology of [21] rests upon the ability to obtain desired predictive distributions and to calculate expectations under these distributions. These authors argue for a cross-validation viewpoint, comparing an observed value against conditional predictive distributions arising from single point deletions. In particular, one of their suggestions is to construct a discrepancy measure (checking function) comparing each observation $Y_{i}^{0}$ against a posterior expectation $E[Y_{i}^{0}\,|\>Y_{\backslash i}^{0}]$ where $Y_{\backslash i}^{0}$ are all but the $i^{th}$ observation inside $Y^{(n)}_{0}$. The conditional expectation after removing the observation $Y_{i}^{0}$ is not too easy to compute and they suggested importance sampling. We suggest one feasible computational strategy based on generative posterior samplers. Using the Bayesian GAN approach [89], one can train a GAN mapping $g(Y_{n-1},X)$ to simulate from $\pi(\theta\,|\>Y_{n-1})$ by passing a random noise variable $X$ through this mapping for a fixed vector $Y_{n-1}$ of length $n-1$. This training is performed solely on fake data. One can then plug a vector $Y_{\backslash i}^{0}$ into $g(Y_{\backslash i}^{0},X)$ to perform iid sampling from the posterior distribution $\pi(\theta\,|\>Y_{\backslash i}^{0})$ for any $i=1,\dots,n$. This sampling is cost-free because it is merely based on passing noise through the trained mapping evaluated at $Y_{\backslash i}^{0}$. The ability to simulate from this posterior makes simulating from $\pi(Y_{i}\,|\>Y_{\backslash i}^{0})$ and, therefore, approximating $E[Y_{i}\,|\>Y_{\backslash i}^{0}]$ feasible. One can, for example, compute the discrepancy measure $Z(Y_{0})=\frac{1}{n}\sum_{i=1}^{n}(Y_{i}^{0}-E[Y_{i}\,|\>Y_{\backslash i}^{0}])^{2},$ which avoids the use of observation $Y_{i}^{0}$ for the predictive check on $Y_{i}^{0}$. A tail area evidence measure can be then obtained by assessing the location of $Z(Y_{0})$ relative to the a reference distribution of $Z(Y)$ where, for example, $Y$ has been simulated from the (posterior predictive under) model $M$.
@@ 5.1 Estimation Consistency
The strength of our generative approach to Bayes factor estimation is that we can evaluate the estimator at any data $Y$, not just the observed data $Y^{(n)}_{0}$. This allows us to recreate distributions of Bayes factor estimators when $Y$ is generated from either model $M_{1}$ and model $M_{2}$. Inspecting entire distributions of Bayes factors $BF_{1,2}(Y)$ provides a more comprehensive image of discriminability among the two models. The quality of our estimator can be gauged by the closeness of the distributions of true and estimated Bayes factors. Figure S20 in the supplement portrays cartoons of plausible shapes of Bayes factor distributions (true and estimated) under the two models. The relative position of the true (or estimated) $BF_{1,2}(Y_{0})$ within the plot provides stronger information than merely deciding whether the Bayes factor is above or below $1$. In addition, we estimate the KL divergence between empirical distributions (kernel density estimates (KDE)). These distances are computed separately under each model and then subsequently weighted by the prior model probability. Similarly, we measure the Spearman’s rho correlation (rank-based) between the estimated and true BF’s for simulated data under each model and take their weighted (by the model prior probability) sum.
We also compute an unbiased estimator of the prior $\pi(M_{j})$ by averaging $\pi(M_{j}\,|\>Y)$ over the simulated data $(Y_{1},...,Y_{T_{0}})^{\prime}\sim P(Y)$, denoted as $\hat{\pi}(M_{j})=\frac{1}{T_{0}}\sum_{i=1}^{T_{0}}\pi(M_{j}\,|\>Y_{i}).$ If $\hat{\pi}(M_{j})$ deviates significantly from true $\pi(M_{j})$, the estimator may not be trusted [72]. In our numerical examples, we assume $\pi(M_{1})=\pi(M_{2})=0.5$.
@@ 5.2 Measures of Surprise
Our simulation-based framework lends itself naturally to inference with Bayes factors based on tail-area surprise measures. While the $p$-value is a tail area (as opposed to likelihood-based) evidence measure, it can be calibrated to obtain a lower bound to a Bayes factor [3]. Before seeing the data, Bayes factors are random variables and we can think of them as a proxy for a test statistic. Evaluating the position of the Bayes factor $BF_{1,2}(Y_{0}^{(n)})$ within the distributions of $BF_{1,2}(Y)$ for $Y\sim\pi(\cdot\,|\>M_{1})$ and $Y\sim\pi(\cdot\,|\>M_{2})$ has a flavor of the prior predictive $p$-value popularized by [13] . We consider surprise measures
@@
which are essentially prior predictive $p$-values using the Bayes factor as a departure statistic. We do not argue that the prior-predictive reference distribution is preferred over alternatives including the conditional predictive $p$-value [3] which is justifiable when there is suspicion that poor priors were used and which is provably superior to various Bayesian $p$-values in large samples [67]. In simulation-based models considered here, however, it might be difficult to come up with appropriate statistics needed for the computation of the conditional predictive $p$-values. While our computational framework enables the use of posterior-predictive $p$-value [30, 68] by simulating $Y$ from the posterior (not prior) predictive under each mode, we focus on the prior predictive $p$-values acknowledging its weaknesses in its prior sensitivity. Using surprise measures for Bayes factors as inferential tools is not a new concept. For example, [87] suggest a simulation-intensive approach to calculate various tail area measures of surprises for Bayes factors and [19] propose a Monte Carlo method to find a decision rule $c$ such that $P(B_{1,2}(Y)\leq c|M_{1})=P(B_{1,2}(Y)\geq c|M_{2})$. Computing such surprise values was also considered in [74] who proposes a Bayes factor design analysis aimed at determining an appropriate sample size $n$ that ensures that with a high probability the sample data from the models exhibit strong Bayes factors in favor of the true model. When closed-form evaluations of Bayes factors are unavailable, the computation of surprise measures can be involved. The ability to rapidly evaluate the Deep Bayes factor mapping for any $Y\in{\rm Supp}(P)$ facilitates the computation of surprise measures. For example, we approximate $p_{1}(\cdot)$ and $p_{2}(\cdot)$ in (13) by replacing $BF_{1,2}(\cdot)$ with an estimator $\widehat{BF}_{1,2}(\cdot)$ and by replacing tail probabilities with empirical proportions based on simulated data $Y$. We denote these approximations by $\hat{p}_{1}(\cdot)$ and $\hat{p}_{2}(\cdot)$. Since the inferential quality of a Bayes factor estimator can be assessed by how well it approximates the tail probabilities in (13), we measure the MSE between the actual surprise and the estimated surprise by $MSE_{surp}=\sum_{j=1}^{2}\pi(M_{j})~{}\hat{\mathbb{E}}_{Y\sim M_{j}}\left[p_{j}(Y)-\hat{p}_{j}(Y)\right]^{2}.$
@@ 5.3 Model Inference
The actual precision in estimating Bayes factors may be less relevant for decision making when there is overwhelming evidence for one of the two models (and the estimation error is appropriately small). It is the quality of decisions with Bayes factor’s that ultimately matters, not the approximation quality of its estimator. Jeffreys’ Bayes factor scale [38] is a suggested interpretation of the evidence strength based on the magnitude of the Bayes factor binned into 10 categories. The coarsest categorization is into two classes, based on the decision boundary at $1$ (values greater than one favor model $1$ and values smaller than one favor model $2$). To quantify the inference quality for various thresholds (not just $1$), we use the area under the ROC curve (AUC) computed as follows. Considering $1$ as a “positive” label and $2$ as “negative”, we can calculate the true positive rate (TPR) and the false positive rate (FPR). Given the decision threshold $c\in(0,\infty),$ we write TPR and FPR as $TPR_{c}(\widehat{BF}_{1,2})=\mathbb{P}(\widehat{BF}_{1,2}(Y)>c|Y\sim M_{1})$ and $FPR_{c}(\widehat{BF}_{1,2})=\mathbb{P}(\widehat{BF}_{1,2}(Y)>c|Y\sim M_{2})$, where $\mathbb{P}$ is the empirical distribution of a simulated sample from the model. As these values depend on the choice of $c$, we can draw the ROC curve representing the relationship between TPR and FPR when $c$ is sent from 0 to $\infty$. The AUC does not penalize biased or improperly scaled estimation.
@@ 6.1 Toy Examples
In this section, we focus on the Negative Binomial vs. Poisson distributions in Example 1. For the other two datasets, see Section S6. This dataset can represent various scenarios with different hyperparameters $(\alpha_{1},\beta_{1},\alpha_{2},\beta_{2})$, which determine the landscape of the decision boundary of the model selection problem. Here, we choose $(\alpha_{1},\beta_{1},\alpha_{2},\beta_{2})=(2,2,4,4)$, which we call Dataset 1 (i), and $(\alpha_{1},\beta_{1},\alpha_{2},\beta_{2})=(1,1,1,1)$ as in Example 1, which we call Dataset 1 (ii). The former is more separable near decision boundaries than the latter in the sense that the probability that $BF_{1,2}(Y)$ is near $1$ is smaller (see, Figure S24 in Section S7). The true Bayes factor is available in a closed form [11]. Letting $Y_{\cdot}=\sum_{i=1}^{n}Y_{i}$,
@@
Because of the explicit likelihood, we can readily use the well-known Bridge sampling [7, 52] and Warp Bridge sampling [51] as baselines. We set the number of MCMC iterations of Bridge and Warp sampling as 150,000 with the burn-in 10,000. We also evaluate the ABC-based Bayes factor estimator [66], which is commonly employed in applications without an explicit likelihood. To compute the ABC-based estimates for a large number of simulated samples within a reasonable timeframe, we adopt a rank-based rule for the ABC algorithm instead of the standard $\epsilon$-acceptance rule. This approach allows for the reuse of ABC samples across different simulated datasets simultaneously, significantly reducing computation time. The detailed algorithm is provided in Section S2, and hyperparameter choices are listed in Table S2 therein. Consequently, the top 120 ABC samples are used to compute the ABC-based estimator in (2) with a small adjustment to avoid numerical explosion. As a result, the range of the ABC-based estimators is approximately [1/120,120]. We vary the sample size from $n=2^{1}$ to $n=2^{7}$, specifically focusing on sample sizes under 1,000. Studying such small sample sizes remains relevant for practical purposes, e.g., in cognitive and psychological science, sample sizes often range from 30 to 60 [72] and a sample size of 180 is considered large [72]. For each $Y_{0}^{(n)}$, we run our algorithm and the baseline methods 20 times to assess estimation stability. As discussed in Section S3, we use both FNNs and BNNs for our algorithm, which we will refer to as DeepBF-FNN and DeepBF-BNN, respectively, where BNNs stand for batch normalization neural networks.
@@
We begin by assessing estimation consistency, as illustrated in Figures 4 and S25 (Section S7) for a sample size $n=128$. The first row of these figures presents scatter plots that compare the estimated $\widehat{BF}_{1,2}(Y)$ with the true Bayes factors $BF_{1,2}(Y)$ from simulated samples $Y$. The scatter plot for DeepBF-BNN suggests unbiased and properly scaled estimates, while DeepBF-FNN exhibits more reserved scaling. In Figure S25, both Bridge and Warp methods yield well-scaled and unbiased estimates. In contrast, Figure 4 shows bias in both methods, although biased estimates may still hold utility within a rank-based decision rule, as reflected by the surprise measures in (13). The ABC method appears non-informative. Additional insights are offered in Figures S27 and S28 (Section S7), which present scatter plots for increasing values of $n$. Here, the ABC method provides more accurate estimations for smaller $n$ values, but these estimations become non-informative as $n$ increases. An increase in the number of ABC samples, currently at 1,200,000, could potentially improve these outcomes. On the contrary, Bridge and Warp show unstable estimates for $Y\sim M_{2}$ but improve as $n$ becomes larger. While DeepBF-FNN tends to outperform DeepBF-BNN, its efficacy wanes with larger $n$, possibly due to the network’s fixed intermediate width across all $n$ values to stabilize optimization (Section S3).
@@
In the second row of Figure 4 and S25 (Section S7), we have empirical distribution plots of the Bayes factors (true and estimated) under the two models. While the KDE shapes of these plots corroborate the scatter plot shapes, we can also see the separability of distributions under each model, and the overall distributional deviation of the estimates. We also visualize the estimation quality for an instance $Y_{0}^{(n)}$ drawn from $M_{2}$. In Figure 4 (f-j), each vertical red line represents the true $BF_{1,2}(Y_{0}^{(n)})$, and the green lines represent its estimates from 20 different trials. For DeepBF-BNN, the green lines cluster closely around the red line, whereas for all other methods, the green lines show more deviation. The blue line is a reference line at the decision threshold of 1. Our algorithm’s estimates tend to be more conservative, closer to the blue line, while the Warp and Bridge estimates are overly optimistic, diverging further from the blue line. The variance in the multiple estimates increases in the following order ABC $<$ Warp and Bridge $<$ DeepBF-BNN $<$ DeepBF-FNN. More comprehensively, the scatter plots among multiple DeepBF-BNN estimates are presented in Figure S29 (Section S7). The plots in Figure S25 (f-j) display similar results, but with Warp and Bridge estimates appearing unbiased and clustering tightly around the true red line.